Three charges, + 35 uC, - 35 uC and + 35 uC are placed at A (0,5cm), B (5cm,0), C(-5cm,0). Calculate the potential energy of the whole system of charges.
Answer #1

qA = 35uC , B = -35 uC, qC = 35 uC

Potential energy U = kq1q2/r

k =9*10^9 N.m^2/C^2

Total potential energy of system U = UAB+UBC+ UAC

rAB = 5*21/2 = 7.07 cm

rAC = 7.07 cm , rBC = 10 cm

U = (9*10^9*35*35*10^-12*10^2)((-1/7.07)+(-1/10)+(1/7.07))

U = - 110.2 J